OpenStudy (anonymous):

HELP 4 MEDAL ! Chris and Justine are purchasing a townhouse and finance $112,900 with a 25-year 6/3 ARM at 4.85% with a 2/12 cap structure. What will their payments be at the beginning of the seventh year assuming they are charged the maximum interest rate for that year? $759.70 $557.09 $674.48 $650.17 I used this formula ->Bm = P[((1+i)^(n*t)-(1-i)^(m))/((1+i)^(n*t)-(1))]

OpenStudy (anonymous):

I plugged in P=112900 I=4.85%/12=0.004041667 N=12 T=25 M=6*12=72 I got 125032 Next I added new info ! I=0.0685/12 = 0.005708333 <--- ( 4.85 + 2 = 6.85%) N=12 T-25-6 = 19 and I got 127231 .. What do I do next .. ?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Hero